OpenStudy (anonymous):

Polygon ABCD is translated to create polygon A′B′C′D′. Point A is located at (1, 5), and point A′ is located at (-2, 1). What is the distance from B to B′ ?

OpenStudy (anonymous):

Well first, you need to see what the translation is, so in this case it is (-3, -4) because A is at (1, 5) but A' is at (-2, 1). 1-3=-2 & 5-4=1 The distance formula is \[d = \sqrt{(x _{2}-x_{1})^{2}+(y_{2}-y _{1})^{2}}\] The distance from B to B' is the same as A to A' or C to C'. So just apply the distance formula to A & you should get 5 as the distance.